Minutes of the 1st Committee Meeting of the Dublin University Rifle Club for A.Y.2001-2

Moderator Steven Watterson Minutes taker Mark Dennehy Present Steven Watterson (Captain),

Benjamin Jones (Treasurer),

Mark Dennehy (Secretary/Match Secretary), Alan Smyth (Chief Range Officer), Conor McDermottroe (Public Relations Officer),

Absent John Keeney (Ammo&Equipment Officer), Jason Sibley, Deirdre Ninaber (Ladies’ Coordinator)

Date 12 June 2001 1800 – 2300 Distribution DURC Committee

Overview of topics

1 Treasurer’s Report 2

2 Committee Position Assignments 2

3 Duties and Responsibilities 2

4 DURC-list 2

5 40th Anniversary Events 3 5.1 Trinity Top Ten . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 5.2 WTSC/DURC/UCDRC Team Event . . . . . . . . . . . . . . . . . . . 3 5.3 RRPC Event . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 5.4 Black Tie Function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 5.5 Trinity College Hartford . . . . . . . . . . . . . . . . . . . . . . . . . . 4 5.6 NSRA/Eley Postal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 5.7 Colours Match . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 5.8 World Student Shooting Games . . . . . . . . . . . . . . . . . . . . . . 4 5.9 Diary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

5.10 Foreign Shooters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 6 Sponsorship 5 7 Trinity News, Trinity Today, University Record 5 8 Membership Drive 5

8.1 Membership fee . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 8.2 Range Tidying . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

8.3 Irish Pony Club . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 9 DURC 50m Open 6 10 Best Newcomers Plate 6 11 NSRA bills 7 12 Summer Opening 7 13 CRO postings 7 14 Ladders 7 15 Squads and Training Practises 7 16 Supervision of Shooters at away competitions 8 17 AOB 9 18 BUSA Top XX 9 19 Three-Position Shooting 9 20 Shotgun Colours 9 21 Social aspect of club 9 22 Playstation 9 23 Bisley 10 24 Request for summer shooting 10 25 Financial dealings rules 10 26 British University Shooting Community 10 27 Security moving to 201 Pearse St. 10 28 Marks Rifles 11 29 Gratutities 11 30 Network Bill 11

1 Treasurer’s Report

Please note that these minutes are confidential to the Club Committee and should not be circulated beyond that group.

1 Treasurer’s Report

Ben reported £280 in the clubs account and that we anticipated a lot of ammunition to be required before the next grant round. Alan agreed saying that since the .22 season had been delayed by F&M that we were in trouble, especially given the amount of ammo we currently have available. Steven suggested that we don’t buy Diablo Sport air rifle pellets again as their quality was far too low, Mark agreed. Steven also suggested that our next .22 order should be 50% Lapua Master and 50% Lapua Multimatch to give better results with some club rifles due to the way these types were manufactured. Ben then asked Mark and Alan to sign the bank form changing singnatories and noted that AIB wanted to see their ID cards for the procedure.

2 Committee Position Assignments

It was decided that Alan would serve as CRO, John as A&E officer (it was decided not to rename the position to Armourer as the mental image of John wielding a broadsword was too funny). The Membership or “Stats” officer would be Ben during the year for ladder preparation but the initial workload of entering several hundred membership records to the database would be split up amongst the committee. At this point there was a discussion on the RCMS project – Mark reported that the basic design was complete and so was the database implementation and that now the tools for inserting and retrieving data and user interaction were being written. The discussion between Conor and Mark sank to technical details a few times but was rescued before the rest of committee fell asleep. Conor was appointed PRO (aka Minister of Propoganda) and it was agreed that Deirdre would be Ladies Coordinator, Steven pointing out that many competitions abroad have seperate ladies teams and that we would need this position. There was a discussion regarding Jasons offical title - Minister without a Portfolio was selected both because Jason is now in final year and because we need people that we can assign to different jobs; but it was decided to ask Jason if he had any individual preference.

3 Duties and Responsibilities

The Match Secretary/PRO interaction was discussed. The PRO will have the job of promoting the club in papers and so forth, as well as taking photographs which was a responsibility that Mark was tired of because of the camera-shyness of most of the club members. The Match Secretary will publish notices of competitions, but the PRO will publicise them on noticeboards further afield than the club. Alan as CRO will be responsible for Booking Grids, this triggered a discussion of the pros and cons of online booking with the RCMS. Mark felt we were speculating without data and suggested a trial period to be held before a decision was made; Conor suggested Hillary Term to avoid losing newbies through technical problems. This was agreed, the length of the period to be decided on at a later date. John as A&E officer was to be responsible for the maintainance of the inventory to prevent a repeat of the delay in producing one for DUCAC this year which caused some levels of friction. It was not felt this would be a heavy task as the inventory does not change much and John could update it either on the fly, or once a term, or after receipt of large consignments of equipment and so forth.

4 DURC-list

5 40th Anniversary Events

Conor pointed out that many posts to durc-internal had been for competitions and would have been better suited to durc-list as well. Mark explained the problems of individuals replying to the durc-list as a whole and how it had been moderated, Conor suggested a restricted posting list of the committee and Mark agreed to look into it and change it to that if possible. Alan noted that his email address might be fouled up for a time as he was changing from an undergrad to a postgrad account, and gave his temporary email address. Task: Adjust durc-list from moderatorship to restricted posting (Mark)

5 40th Anniversary Events

Mark gave a brief review of the dates relating to the club: the modern club was founded 40 years ago, the original club is 163 years old and the earliest documented match took place in college 161 years ago.

5.1 Trinity Top Ten

Mark reported that we had had one entry from DRC, interest from Trinity College Hartford and from clubs in Austrailia. Steven agreed to talk to clubs from the Isle of Man (as he was from there) and Edinburgh (where he had shot as an undergrad), as well as other scottish clubs. Mark also gave an overview of the changes regarding the introduction of two new leagues in the Top Ten. Task: Talk to clubs in Isle of Man and Scotland about entering Trinity Top Ten (Steven)

5.2 WTSC/DURC/UCDRC Team Event

This was introduced by Mark as being an event Geoff Cooney of WTSC wanted to set up. At this point Mark mentioned the new Air Rifle Classification System that Geoff had been working on as Air Rifle Coordinator in the NTSA and explained the rough workings of it. Steven was surprised that he didn’t make the top grade under the new system. Concern was expressed over the larger statistics workload; Mark agreed but said PCs and spreadsheets made this not so onerous. Concern was also expressed over the increase in prizes; Mark said prizes were not given out for each sub-grade, but only for the four classes.

It was agreed that Mark, Geoff and the UCD Match Sec would meet and organise this event, which was to be listed in the DURC calendar as a Celebratory event to keep in theme with the 40th Anniversary. Task: Organise DURC/WTSC/UCDRC event in concert with Geoff Cooney & UCDRC Match Secretary (Mark)

5.3 RRPC Event

Mark pointed out that RRPC also celebrate their 40th anniversary this year and that they had a 40th anniversary shoot planned. He suggested reviving the RRPC vs. DURC matches which were last shot in 1993 and whose shield was still on the range. This was agreed, the details to be arranged between Mark and the RRPC Match Sec. Task: Organise RRPC/DURC event in concert with RRPC Match Secretary (Mark)

5.4 Black Tie Function

Mark explained what he meant by this event, a formal dinner inviting the historical members of DURC in the TCD Dining Hall. He stated that we should try to find the founding members current status and invite them if possible. It was stated that this would not be a general shooting event, but a formal dinner. DUFC and DUEC committees would be invited as the club has a long association with these clubs. Ben pointed out that the clubs funds were low and that the function could not put a large load on the clubs finances. Steven wondered about the order of magnitude of the attendance, 100 was agreed on as a good estimate. Mark agreed to talk to Trevor West about the function to secure his assistance. Alan said all that was needed was effectively a second sitting of Commons. Ben pointed out at this point that Cyril, DUCAC treasurer, was due to be promoted and no replacement had yet been named. Steven said we should start moving on the function in November, and explained how the rifle club in Edinburgh handled its annual ball, which sounded like a very similar event. Steven was concerned over the marketability of the event – Mark said that we would be inviting people to a formal black-tie function in the Trinity College Dining Hall, which was prestigous enough to nullify the problem. Conor and Ben agreed that pricing tickets could not approach Trinity Ball levels. Mark related the levels for the RO dinner, and Ben talked about the pricing of same. Ben also pointed out the amount the club had spent on beer last year was far too high (£600 approximately) Task: Talk to Trevor West about Black Tie event to secure his assistance (Mark)

5.5 Trinity College Hartford

Steven wondered if anyone had heard of Hartford before, to see what its academic standing was like. Conor had heard of them but had not heard opinions on their standing. There followed an explaination of correspondance with their captain by Mark and a discussion of possible formats for a postal competition as TCH were DURCs first opponents outside the British Isles in 1967 - when we won. Disciplines, ranges were discussed. There was confusion over TCHs discipline, Mark thought perhaps it was a misspelling in their email that they shot 50ft, when they meant 3P 50metres, and he agreed to follow up on that. Steven noted that there was a case for comparing scores from 15yd indoor to 25yd indoor if necessary. Task: Follow up on TCH event (Mark)

5.6 NSRA/Eley Postal

Mark related the faxes and emails he had sent to the NSRA regarding the second round of cards that had been submitted and their lack of response. Steven agreed to follow up as he had contact details for the NSRA. Task: Contact NSRA regarding NSRA/Eley Postal cards (Steven)

5.7 Colours Match

Mark reported that the Colours was now finally offically recognised by CUSAI and that we had been told we should enter the Intervarsity Event of the Year award next year and thus we would be eligible to attend Sporting Commons.

5.8 World Student Shooting Games

Mark reported that this, the first time shooting was in the world student games, was being organised for this year, to be shot in the Czech Republic in September 2002; also that Dave Cooney was enthuastic about entering a team from UCDRC and that Pat O’Brien was attempting to locate details on the event for the NTSA. It was felt we should enter, though we would probably not do well - but it was pointed out that as UCDRC would enter that we would at worst only come second last. Mark wondered if postgrads would be elegible to enter, and agreed to follow up on enquiries. Task: Follow up on gathering information re World Student Shooting Games (Mark)

5.9 Diary

Steven noted that a good diary was needed to keep track of when entrys and so forth were required for these competitions. Mark agreed. Task: Maintain diary with dates for competition entrys and so forth (Mark)

6 Sponsorship

5.10 Foreign Shooters

Ben wondered if we could get English shooters to attend competitions in Ireland. Steven said it was a good idea but hard to do as not many Universities in England shot Air Rifle, and Mark pointed out the problem with the fee for the temporary Gun licence in the Republic. Mark related the details of how the Air Nationals had circumvented this by loaning actions to english competitiors, who brought their own stocks, Steven was very impressed by this idea. Mark suggested that an event where we shot against english shooters should be held in the North where temporary licences were free; Ben pointed out that DUCAC would not fund this.

6 Sponsorship

Mark gave a brief outline of the efforts made to date on this front with the letter to Eircom to which no reply had been received. It was felt that the amount (£5000) asked for had been to high, Mark stated that it was not as he had seen the cash flow levels within Eircom and that they could easily support such an outlay while other companies might not. Steven said that applying to shooting manufacturers was a better policy and that we should observe their percieved market penetration when considering applying. Ben suggested banks and other institutions considering advertising at students, such as Meteor and other mobile phone companies. Steven suggested asking Eley and so forth for donated ammunition - but that proved less useful than expected because of how DUCAC funds our ammunition. Mark suggested that if they would cover prizes only, it would be a great help as these are not funded by DUCAC. Ben suggested a greater public image would be beneficial to attracting sponsorship which was agreed. Steven will agree to apply to the Bank of Ireland and Allied Irish Bank for sponsorship, Mark to forward him the L^{A}TEX files he used for the Eircom letter. Task: Apply for sponsorship to AIB, BoI (Steven) Task: Pass on L^{A}TEX files to Steven from Eircom bid (Mark)

7 Trinity News, Trinity Today, University Record

It was decided the PRO would prepare articles for these college publications, including a 40th anniversary article for the Trinity News. Alan mentioned that he had a contact in the University Record which would give us priority in access to publication there. Mark was to pass on the photographs he had on file for DURC to Conor. Task: Prepare articles for college press publications (Conor)

8 Membership Drive

Ben explained DUCACs proposed second week of Freshers demos in Luce Hall. This was seen as a good idea by all. Steven queried the amount of time and scheduling for demos. Steven also said a list was needed of those we could nab for RO duty during Fresher’s Fortnight. There was a query as to where Deidre hailed from so we’d know if she was available for the fortnight.

Steven then introduced the idea of running participation demos on the range in fresher’s week, so that newly joined-up members could immediately get to try air rifle for ten shots or so. Alan agreed, saying that opening in the first week of term was vital as in previous years the initial enthuasism had been lost by the time we opened in the second week. The general feeling was that starting quickly was an excellent idea.

Mark brought up the leaflets, which were acknowleged as having been a good idea, and the difficulties in printing the quantities required. Printing in small batches over a long time was forwarded as a better idea, but commercial printing was suggested by Mark as the best idea, giving a panel of the leaflet over to advertising for a commercial

9 DURC 50m Open

sponsor like Reads or Meteor or whomever. Conor suggested leaving these leaflets in the packs given out to the Freshers on their enrollment.

8.1 Membership fee

It had been proposed to change from £2 to £5, of which £2 would be returned in the form of range credit: but it was felt that this would reduce membership to nothing. Ben pointed out that £1 was give to DUCAC for every member and that the end-of-week party was paid for by the other £1 so that we couldn’t reduce our fee, and that societies didn’t have to pay this levy, only sports clubs. The idea of changing the fee was rejected, and it was decided to retry asking for sponsorship from beer companies or O’Briens or Red Bull or any such company.

8.2 Range Tidying

Alan announced he would “tidy the range through delegation” Task: Tidy range (Alan)

8.3 Irish Pony Club

Mark related the details of the IPCRA report at the NRPAI agm and said that we should be targeting IPC shooters for DURC as their standard was very high. He agreed to talk to the DUEC committee further on this. Conor suggested a joint DURC/DUEC membership offer, this was seen as a good idea, but that it might bring financial problems, though if sponsorship for the end-of-week party could be arranged, that problem would solve itself. Task: Meet with DUEC committee regarding IPC member recruitment (Mark)

At this point a 15 minute recess was called so that the captain could check the cricket scores, the treasurer could check on his physics experiment and the secretary could collect some paperwork and refreshments. Alan left at this point and all reconvened after England had been beaten by Pakistan.

9 DURC 50m Open

Steven suggested the end of August for a date; Mark warned about checking for conflicts with other events as the 50m calendar had been badly affected by the F&M crisis and many events had backed up. The MRC venue was discussed as an alternative, and was felt to be an acceptable venue but the problem was the amount needed for range hire, as FSC was felt to be free given our contribution the year earlier. It was agreed that Mark would ask John to secure more information on this as it was John’s local range. Mark was also to secure accurate information on the NTSA levy amount. There was a discussion on the amounts charged as fees but nothing was decided. The prizes were also discussed, with Steven feeling that in the higher grades, silverware prizes were commonplace enough to be not an effective attractor and that alternatives should be considered. Wine was suggested, as was rifle equipment - but that was dismissed as shooters at that level had all their own equipment. Task: Find out about NTSA levy amounts (Mark) Task: Gather information about MRC, including range hire fees (John)

10 Best Newcomers Plate

Conor inquired about his plate; Mark replied that he had been unable to locate Anna Hickeys current location but was still trying. Task: Locate Anna Hickey and secure return of plate (Mark)

11 NSRA bills

11 NSRA bills

Ben said that the NSRA had not yet sent the books or targets that we’d ordered; Steven gave Ben the invoice he had received from the NSRA for range plans. Ben agreed to follow up on the NSRA order. Task: Follow up on NSRA order for targets and books (Ben)

12 Summer Opening

It was pointed out that there had been enquiries about shooting during the summer. Mark suggested that those that shot then should be required to enter the Trinity Top Ten; it was decided that it should not be mandatory but that it should be strongly advocated by the ROs. Mark was also annoyed at his friends constantly choosing dates for getting married when a competition he wanted to attend was on. It was decided after debate that the range should open one or two nights a week on an informal basis, but that those wishing to shoot had to phone the club or email rifle.club@tcd.ie to say so: Mark was to arrange with ISS to redirect that address to the committee list.

Ben noted that Aisling would not be around during the summer but would be there for the Freshers week. Task: Arrange for rifle.club@tcd.ie to be redirected to committee list (Mark)

13 CRO postings

It was agreed that the CRO would be formally responsible for posting weekly notices to durc-internal with the weeks notes, as Mark had done on a few occasions last year.

14 Ladders

The system of ladders in the club was discussed. The idea of a challenge-based ladder was considered and rejected on the grounds that it would be akward to maintain and that past experience with the system showed that enthuasiasm waned very quickly. There followed a discussion on possible alternatives, including Ben’s production at the AGM of a last-12 ladder, and the idea of a rapidly changing ladder was seen as an good idea, but the question of which ladder got the prize was a problem. Conors ladder from last year was also pointed out, as his early scores had unfairly handicapped him at the end of the year, while Adrian Costigan had beaten him on the overall ladder because he had not shot as many cards. After discussion of technical implementation and Conor and Mark assuring the committee that it would not be a major problem to implement, it was decided that this year two ladders would be maintained by Ben using the RCMS; a weekly ladder carrying the best 5 of the last 6 cards shot, and the yearly ladder which would be the average of the top 50% of the scores that shooter had put in. The top novice in the weekly ladder would pay RO rates on Winchester ammunition or Air Pellets for that week; presumably the top experienced shooter would as well if he/she were not an RO.

Steven pointed out that as money was now involved, the updating of the ladders would have to be more formalised; Ben said he had done so on a weekly basis last year; it was agreed that this was sufficent, with the addition that should Ben be away for a week at conference or other, that he would appoint a deputy for that period. Task: Write chunks of RCMS to handle ladders as discussed (Conor,Mark)

15 Squads and Training Practises

Mark felt that the squads had been a good idea last year but not sufficently backed up to give excellent results on a large scale. Steven felt that the arrangements for selection needed to be less formal than a qualification score. The question of coaching was raised,

16 Supervision of Shooters at away competitions

specifically where to find coaches as Steven, Mark, John, Ray and Keith were felt to be the only people in the club technically adept enough to coach. Mark agreed to ask members of the National Squad to come down to the range and coach for a night; Ben said we could pay them £15 per night, which was felt to be a pittance.

Steven felt that selection for the squad needed to be less formal, and that it should be more on the basis of motivation as opposed to technical skills. He was hoping for a situation where we could specify what the squad would train in: Mark pointed out that some people disliked some disciplines enough to avoid them and that we should instead what we would be training in and let the squad attend if they did that discipline, as per the National Squad’s arrangements. Task: Invite members of National Squad to train DURC Squad (Mark)

16 Supervision of Shooters at away competitions

Mark described the problems that had been encountered at the Nationals with ammunition distribution, firing point allocation, and other technical issues. He read the minutes of the committee lunch where it had been suggested that a “nanny” system should be used, where one person (with a deputy to take over when the nanny was shooting) would be responsible for organising details, issuing and tracking ammunition and so forth. Steven suggested instead a “buddy” system, whereby each shooter would be paired with another. The more experienced shooter would then be assisted by the junior, who would learn by osmosis. This had the advantage of one-to-one supervision, as well as being of real benefit to both shooters as setting up a firing point in RRPC was an onerous task to do in three minutes. This was felt by all to be an excellent suggestion and it was agreed to implement it immediately for trial this weekend at the FSC Open.

The problems posed by Ger Ryans behaviour at the Nationals were then brought up. It was felt by Mark and those who had attended that his behaviour reflected poorly on the club, and that it was not nervousness due to the continuation of this behaviour at the AGM. A written Code of Conduct for the clubs away trips was suggested and accepted as a worthwhile measure. Steven agreed to prepare on with the NTSA’s code as a basis, which Mark agreed to locate and pass on. The Code was to be on a single A4 sheet which would be signed by the competitor before going to the competition. The timing of the signing was debated and the agreement was that the competitor should sign before going, but that we would allways have a copy handy for unexpected arrivals at competition.

Further to this, it was agreed that from now on, attendance at competitions was to be initiated by an invitation from the ROs with guidelines issued for them, but ultimately at their discresion. A grid-based system for attending was agreed as the best system to continue to use. Task: Prepare a Code of Conduct document for the club along lines of NTSA code (Steven) Task: Supply Steven with NTSA code of conduct (Mark)

17 AOB

17 AOB

Steven had several other points to add to the meeting.

18 BUSA Top XX

Steven suggested entering the BUSA Top XX team competition and the Top XX individual competitions this year. This was agreed as an excellent idea and Steven agreed to get the contact details and Mark would follow up. Task: Get contact details for BUSA (Steven) Task: Follow up (Mark)

19 Three-Position Shooting

Mark had been suggesting for a year that we encourage 3P shooting in the club. Steven said he had a problem with that in regards to safety as the change in shooting rythym might cause someone to set off a round with the gun vertical in standing or kneeling position. Mark pointed out that this had happened last year from the Prone position, but agreed it was a problem. Ben suggested that we reintroduce 3P Air Rifle as a discipline to counter these problems so that we could train for 3P in the club. Steven and Mark agreed this was an excellent idea, not just as a compromise but as a discipline in its own right. Mark asked if it would get a night per week, after some debate it was agreed Conors suggestion of switching to 3P Air after 2000 on an Air night would be best. The idea of a 3P Air colours against UCD was brought up - Mark will bring up the idea with UCDRC and see how they feel. Steven pointed out that at the point at which you shoot from for prone and kneeling with air, you require a platform, as you are too low with respect to the target. This was seen as a minor problem, and an excellent project for Keith White. Task: Discuss 3P Air colours with UCDRC (Mark)

20 Shotgun Colours

The shotgun colours was brought up at this point - Steven felt that while it was enjoyable as a novelty competition, it worried him that we would be holding a competition in a discipline we did not train for. The social aspect of the competition was put forward as sufficent cause for the event. No decision was reached.

21 Social aspect of club

Steven proposed a general rule of thumb for this year, that the ROs tell members on the night which pub they would attend after shooting and when so that they could get together in an informal setting for a chat - this to increase the members interest in the club. This was seen as a good idea, though many of the ROs would not drink that often, and that timing was a problem as the shooting ran on so late. No final decision was reached, though it was felt to be a good idea.

22 Playstation

Steven is concerned over the effect the PC in the lounge was having on the social aspect of the club as many people came in and read email or played games and never talked to the others. Mark agreed, pointing out that the PC was never intended for this purpose and that the current webserver was to have been replaced by it. Steven did not know, however, what to do with the PC, so it was agreed to consider it for a later discussion.

23 Bisley

23 Bisley

Steven worried that we had not been entered for Bisley Week. Mark pointed out the confusion over the trip and whether it had been one member going on his own or a group representing the club. He had therefore not been asked to look at it until just before he was running off to the Isle of Man and had not arranged anything. Mark agreed to check the entry status with the NSRA. Task: Check status of entrys to Bisley Week with NSRA (Mark)

24 Request for summer shooting

A request for a staff member to shoot during the summer was considered. Noone had any problems with it as there would be .22 training anyway this summer. It was agreed that Mark would post a notice to durc-list explaining the informal arrangements for summer shooting and that it was also not a problem for Paul Dunne to shoot. Conor said he was able to RO most afternoons if needs be. Ben asked if an RO duty roster would be needed, it was agreed that if it was, Alan would arrange it. Task: Post notice to durc-list regarding details of summer shooting (Mark)

25 Financial dealings rules

Ben agreed to produce a one-page letter detailing the rules for dealing with DURC for finances. This was felt to be necessary as the NTSA and other clubs were not providing receipts or invoices which was causing problems as DUCAC insisted on these. Task: Produce list of procedures for clubs dealing financially with DURC (Ben)

26 British University Shooting Community

Steven proposed that we get involved in the British Universities shooting scene, as the scores in the club have dropped off over the years and we are becoming slightly complacent. Mark thought this was an excellent idea. Ben pointed out that DUCAC would not fund UK trips. Mark then showed the Competition Attendance Reports he was giving to DUCAC to increase their perception of the club. Steven suggested a subsidised trip as opposed to a fully funded one, there was a discussion on the fee’s level, during which the idea of means testing was rejected, and it was agreed that £100 was too high and £50 was acceptable but might not cover costs. Steven had suggested attending the Scottish Universities Indoor Rifle championships in Perth, which is 3 cards at 25yds followed by a social event, with teams of 8. It was felt that a team event between universities was a good case for DUCAC to provide some funding. It was suggested Ben personally inquire about the levels of UK funding given to other clubs by contacting their treasurers, where these were personally known. Task: Contact other treasurers about UK funding levels (Ben)

27 Security moving to 201 Pearse St.

Mark recounted his meeting with Pat Moray regarding the move of the security offices to 201 Pearse St. While the Front Arch office would remain open, it was not felt acceptable by the committee that the rifles and ammunition would then be supervised by the Accomodations office by day. Pat was also not happy at the current arrangement where ammunition is stored beside rifles. It was agreed that Mark would meet Pat again and clarify whether 201 had space for Ammunition and we would purchase a new gun safe for the range to store the remaining rifles on the range. Mark would also clarify where the keys to the range would be stored - it was not felt best to store them in Front Arch. Task: Meet Pat Moray again and clarify storage situation for Ammo and Keys to range in 201 Pearse St. (Mark)

28 Marks Rifles

28 Marks Rifles

Mark having bought two rifles (a 2013 and a 2002) in Munich the previous week, asked for offical permission to store them in the club (having asked informally before buying them). This was agreed without debate.

29 Gratutities

Mark proposed a formal bribery fund be allocted in the Treasurers records to fund christmas cards, birthday cards and small gifts to relevant people to smooth the running of the club. This was agreed as a good idea.

30 Network Bill

ISS not having sent the remainder of the bill for the club’s network hire, Mark agreed to remind them to do so. Task: Ask ISS for bill for network hire (Mark)

T.B.N.O.B.T.M.W.D.C.

Steven Watterson, Club Captain

Mark Dennehy, Club Secretary

30 Network Bill

Tasks agreed on at meeting

Mark Adjust durc-list from moderatorship to restricted posting (12 June

2001) [/222] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . /2

Steven Talk to clubs in Isle of Man and Scotland about entering Trinity

Top Ten (12 June 2001) [/242] . . . . . . . . . . . . . . . . . . . . . . /2

Mark Organise DURC/WTSC/UCDRC event in concert with Geoff

Cooney & UCDRC Match Secretary (12 June 2001) [/264] . . . . . . . /2

Mark Organise RRPC/DURC event in concert with RRPC Match Sec-

retary (12 June 2001) [/275] . . . . . . . . . . . . . . . . . . . . . . . . /2

Mark Talk to Trevor West about Black Tie event to secure his assistance

(12 June 2001) [/309] . . . . . . . . . . . . . . . . . . . . . . . . . . . . /2

Mark Follow up on TCH event (12 June 2001) [/327] . . . . . . . . . . . . /2 Steven Contact NSRA regarding NSRA/Eley Postal cards (12 June 2001) [/336] /2 Mark Follow up on gathering information re World Student Shooting

Games (12 June 2001) [/359] . . . . . . . . . . . . . . . . . . . . . . . /2

Mark Maintain diary with dates for competition entrys and so forth (12

June 2001) [/368] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . /2

Steven Apply for sponsorship to AIB, BoI (12 June 2001) [/409] . . . . . . /2 Mark Pass on L^{A}TEX files to Steven from Eircom bid (12 June 2001) [/410] /2 Conor Prepare articles for college press publications (12 June 2001) [/421] . /2 Alan Tidy range (12 June 2001) [/469] . . . . . . . . . . . . . . . . . . . . /2 Mark Meet with DUEC committee regarding IPC member recruitment

(12 June 2001) [/483] . . . . . . . . . . . . . . . . . . . . . . . . . . . . /2

Mark Find out about NTSA levy amounts (12 June 2001) [/516] . . . . . /2 John Gather information about MRC, including range hire fees (12 June

2001) [/518] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . /2

Mark Locate Anna Hickey and secure return of plate (12 June 2001) [/526] /2 Ben Follow up on NSRA order for targets and books (12 June 2001) [/535] /2 Mark Arrange for rifle.club@tcd.ie to be redirected to committee list (12

June 2001) [/556] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . /2

Conor,Mark Write chunks of RCMS to handle ladders as discussed (12 June

2001) [/597] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . /2

Mark Invite members of National Squad to train DURC Squad (12 June

2001) [/622] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . /2

Steven Prepare a Code of Conduct document for the club along lines of

NTSA code (12 June 2001) [/665] . . . . . . . . . . . . . . . . . . . . . /2

Mark Supply Steven with NTSA code of conduct (12 June 2001) [/666] . . /2 Steven Get contact details for BUSA (12 June 2001) [/680] . . . . . . . . . /2 Mark Follow up (12 June 2001) [/681] . . . . . . . . . . . . . . . . . . . . /2 Mark Discuss 3P Air colours with UCDRC (12 June 2001) [/705] . . . . . /2 Mark Check status of entrys to Bisley Week with NSRA (12 June 2001) [/748] /2 Mark Post notice to durc-list regarding details of summer shooting (12

June 2001) [/763] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . /2

Ben Produce list of procedures for clubs dealing financially with DURC

(12 June 2001) [/774] . . . . . . . . . . . . . . . . . . . . . . . . . . . . /2

Ben Contact other treasurers about UK funding levels (12 June 2001) [/799] /2 Mark Meet Pat Moray again and clarify storage situation for Ammo and

Keys to range in 201 Pearse St. (12 June 2001) [/817] . . . . . . . . . /2

30 Network Bill

Mark Ask ISS for bill for network hire (12 June 2001) [/841] . . . . . . . . /2